# Dual-Objective Reinforcement Learning with Novel Hamilton-Jacobi-Bellman Formulations

### v1

@@ Abstract
In this work, we extend recent advances that connect Hamilton-Jacobi (HJ) equations with RL to propose two novel value functions for dual-objective satisfaction. Namely, we address: (1) the Reach-Always-Avoid problem – of achieving distinct reward and penalty thresholds – and (2) the Reach-Reach problem – of achieving thresholds of two distinct rewards. In contrast with constrained Markov processes or temporal logic approaches, we are able to derive explicit, tractable Bellman forms in this context by decomposing the problems into reach, avoid, and reach-avoid problems to leverage the recent advances. Moreover, we leverage our analysis to propose a variation of Proximal Policy Optimization, dubbed (DO-HJ-PPO), to solve this class of problems and demonstrate that it bests other baselines in safe-arrival and multi-target tasks, providing a new perspective on constrained decision-making.
I Related Works
Many in learning and autonomy have considered balancing safety and liveness. A few particularly relevant topics are mentioned here. Constrained Markov Decision Processes (CMDPs) are a popular approach to transform constraints into a Lagrangian [1, 2, 3, 4, 5], however, this often requires intricate reward engineering and parameter tuning to balance the combined objective. Similarly, Multi-Objective RL solves the Pareto optimal solution of vector-valued rewards but is not focused on priority-scalarized problems [6, 7, 8]. Goal-Conditioned RL [9, 10, 11] and Linear/Signal Temporal Logic RL [12, 13, 14, 15, 16, 17] generally learn to solve multiple tasks at once, by augmenting the problem to a surrogate problem or automaton, but this is a notoriously challenging approach and there are often no guarantees for the relation between the surrogate and original problem. In this work we are able to derive explicit forms for dual-objective Bellman equations that yield the optimal policy by augmenting the state. This leads to direct approaches for learning the dual-objective values, which is proven to yield improved performance. To do this, we build on traditional dynamic-programming methods that solve Hamilton-Jacobi-Bellman (HJB) equations, and specifically those that connect HJB and RL theories [18, 19, 20, 21].
II Problem Definitions
@@
Consider a Markov decision process (MDP) $\mathcal{M}=\langle\mathcal{{S}},\mathcal{{A}},f\rangle$ consisting of finite state and action spaces $\mathcal{{S}}$ and $\mathcal{{A}}$, and unknown dynamics $f$ that define the deterministic transition $s_{t+1}=f(s_{t},a_{t})$. Let an agent interact with the MDP by selecting an action with policy $\pi:\mathcal{{S}}\to\mathcal{{A}}$ to yield a state trajectory $s_{t}^{\pi}$, i.e. $s_{t+1}^{\pi}=f\left(s_{t}^{\pi},\pi\left(s_{t}^{\pi}\right)\right).$
@@
In this work, we consider the Reach-Always-Avoid (RAA) and Reach-Reach (RR) problems, which both involve balancing two objectives, which are each specified in terms involving a best reward or worst penalty encountered over time. For the RAA problem, let $r,p:\mathcal{{S}}\to\mathbb{R}$ be a reward to be maximized and a penalty to be minimized. For the RR problem, let $r_{1},r_{2}:\mathcal{{S}}\to\mathbb{R}$ be two distinct rewards to be maximized. The agent’s overall objective is to maximize the worst-case outcome between the best-over-time reward and worst-over-time penalty (RAA) or the two best-over-time rewards (RR):
[…]
where $q=-p$. Note that we have introduced this negated penalty $q$ in order to establish a "higher-is-better" convention across subtasks as to compose them in an "and" fashion via the $\min$ operation. For conceptual ease, we recommend the reader still think of trying to minimize the largest-over-time penalty $p$ rather than maximize the smallest-over-time $q$.
Figure 1: DDQN Demonstration of the RAA & RR Problems We compare our novel formulations with previous HJ-RL formulations (RA & R) in a simple grid-world problem with Double-Deep $Q$ Learning. The hazards are highlighted in red, the goal in blue, and trajectories in black (starting at the dot). In both models, the agents actions are limited to left and right or straight and the system flows upwards over time.
@@
As the problem names suggest, these optimization problems are inspired by (but not limited to) tasks involving goal reaching and hazard avoidance. While these problems are thematically distinct, they are mathematically complementary. The respective values corresponding to a policy $\pi$ are
$V_{\textup{RAA}}^{\pi}(s)$ $=\min\left\{\max_{t}r(s_{t}^{\pi}),\>\min_{t}q(s_{t}^{\pi})\right\}$
$V_{\textup{RR}}^{\pi}(s)$ $=\min\left\{\max_{t}r_{1}(s_{t}^{\pi}),\>\max_{t}r_{2}(s_{t}^{\pi})\right\}.$
Observe that the minimums and maximums over time make these values fundamentally different from more typical "sum-of-discounted-reward" objectives considered in RL [22]. Moreover, while both returns involve two objectives, these objectives are combined in worst-case fashion to ensure dual-satisfaction.
III Reachability and Avoidability in RL
Prior works [18, 19] study the reach $V_{\textup{R}}^{\pi}$, avoid $V_{\textup{A}}^{\pi}$, and reach-avoid $V_{\textup{RA}}^{\pi}$ values, respectively defined by
$V_{\textup{R}}^{\pi}(s)$ $=\max_{t}r(s_{t}^{\pi}),$
$V_{\textup{A}}^{\pi}(s)$ $=\min_{t}q(s_{t}^{\pi}),$
$V_{\textup{RA}}^{\pi}(s)$ $=\max_{t}\min\left\{r(s_{t}^{\pi}),\min_{\tau\leq t}q(s_{\tau}^{\pi})\right\},$
@@ III Reachability and Avoidability in RL
for which special Bellman equations [18]. To put these value functions in context, assume the goal $\mathcal{G}$ is the set of states for which $r(s)$ is positive and the hazard $\mathcal{H}$ is the set of states for which $q(s)$ is non-positive. Then $V_{\textup{R}}^{\pi}$, $V_{\textup{A}}^{\pi}$, and $V_{\textup{RA}}^{\pi}$ are positive if and only if $\pi$ causes the agent to eventually reach $\mathcal{G}$, to always avoid $\mathcal{H}$, and to reach $\mathcal{G}$ without hitting $\mathcal{H}$ prior to the reach time, respectively. The Reach-Avoid Bellman Equation (RABE), for example, takes the form [19]
@@
and is associated with optimal policy $\pi_{\text{RA}}^{*}(s)$ (without the need for state augmentation, see Section A in the Supplementary Material). This formulation does not naturally induce a contraction, but may be discounted to induce contraction as in [19] by (implicitly) defining $V_{\textup{RA}}^{\gamma}(z)$ for each $\gamma\in[0,1)$ via
$V_{\textup{RA}}^{\gamma}(s)$ $=(1-\gamma)\min\{r(s),q(s)\}\quad+\quad$
$\gamma\min\left\{\max\left\{\max_{a\in\mathcal{{A}}}V_{\textup{RA}}^{\gamma}\left(f(s,a)\right),r(s)\right\},q(s)\right\}.$
@@
These prior value functions and corresponding Bellman equations have proven powerful for these simple reach/avoid/reach-avoid problem formulations. In this work, we generalize these results to the aforementioned broader class of problems.
IV The need for augmenting states
@@
Unlike for the reach/avoid/reach-aviod problems, the goal of choosing a policy $\pi:\mathcal{{S}}\to\mathcal{{A}}$ is inherently flawed for the RR and RAA problems without state augmentation. In considering multiple objectives over an infinite horizon, situations arise in which the optimal action depends on more than the current state, but rather the history the trajectory. An example clarifying the issue is shown in Figure 2.
@@
Figure 2: Examples where a Non-Augmented Policy is Flawed In both MDPs, consider an agent with no memory. (Left) For a deterministic policy based on the current state, the agent can only achieve one target (RR), as the policy must associate the middle state with either of the two possible actions. (Right) In the RAA case, assume the robot must avoid the fire at all costs and would prefer to not encounter the peel, but will do so if needed. The optimal decision for the current state depends on state history, specifically on whether the robot has already reached the target state or not.
@@
To allow the agent to use relevant aspects of its history, we will henceforth consider an augmentation of the MDP with auxiliary variables. A theoretical result in the next section states that this choice of augmentation is sufficient in that no additional information (nor even the use of a stochastic policy) will be able to improve performance under the optimal (deterministic) policy.
IV-A Augmentations

### v2

@@ Abstract
Hard constraints in reinforcement learning (RL) often degrade policy performance. Lagrangian methods offer a way to blend objectives with constraints, but require intricate reward engineering and parameter tuning. In this work, we extend recent advances that connect Hamilton-Jacobi (HJ) equations with RL to propose two novel value functions for dual-objective satisfaction. Namely, we address: 1) the Reach-Always-Avoid (RAA) problem – of achieving distinct reward and penalty thresholds – and 2) the Reach-Reach (RR) problem – of achieving thresholds of two distinct rewards. In contrast with temporal logic approaches, which typically involve representing an automaton, we derive explicit, tractable Bellman forms in this context via decomposition. Specifically, we prove that the RAA and RR problems may be rewritten as compositions of previously studied HJ-RL problems. We leverage our analysis to propose a variation of Proximal Policy Optimization (DOHJ-PPO), and demonstrate that it produces distinct behaviors from previous approaches, out-competing a number of baselines in success, safety and speed across a range of tasks for safe-arrival and multi-target achievement.
1 Introduction
The development of special Bellman equations from the Hamilton-Jacobi (HJ) perspective of dynamic programming (DP) has illustrated a novel route to safety and target-achievement in reinforcement learning (RL) Fisac et al. (2019); Hsu et al. (2021). In comparison with the canonical RL discounted-sum cost and corresponding additive DP update, these equations, namely the Safety Bellman Equation (SBE) and Reach-Avoid Bellman Equation (RABE), propagate the minimum (worst) penalty and maximum (best) reward, yielding a value function defined by the outlying performance of a trajectory. In mission-critical applications, where avoiding failure is a necessary condition, these equations have proved invaluable in the field of safe control Mitchell et al. (2005); Ames et al. (2016). By focusing on extremal values rather than discounted sums, the HJ-RL equations induce behaviors that act with respect to the best or worst outcomes in time-optimal fashions, performing far more safely than Lagrangian methods Ganai et al. (2023); So et al. (2024). Accordingly, these updates yield policies with significantly improved performance in target-achievement and obstacle-avoidance tasks over long horizons Yu et al. (2022a; b), relevant to fundamental and practical problems in many domains.
Figure 1: Depiction of the Reach-Always-Avoid (RAA) and Reach-Reach (RR) Tasks. In the RAA tasks, the zero-level set of the rewards (goals) and penalties (obstacles) are depicted in green and red respectively, while in the RR problem, the zero-level set of the two rewards (two goals) are depicted in green and blue. The RAA value is defined by the minimum of the minimum penalty and maximum reward, inducing the agents to enter the goals at some time without ever entering the obstacles. The RR value is defined by the minimum of the two maximum rewards, inducing the agents to enter both goals at some time.
In this work, we advance the existing HJ-RL formulations by generalizing them to compositional problems. To date, the HJ-RL Bellman equations are limited to three operations: Reach (R), wherein the agent seeks to reach a goal (achieve a reward threshold), Avoid (A), wherein the agent seeks to avoid an obstacle (avoid a penalty threshold), and Reach-Avoid (RA), where the agent reaches a goal while avoiding obstacles on the way. In this light, we extend the HJ-RL Bellman equations to two complementary problems concerned with dual-satisfaction, namely the Reach-Reach (RR) problem for reaching two goals and the Reach-Always-Avoid (RAA) problem for continuing to avoid hazards after reaching a goal, demonstrated in Figure 1. We prove that the RAA and RR have a fundamental structure such that their Bellman equations may be decomposed into combinations of SBEs and RABEs. From this theory, we devise DOHJ-PPO, a novel algorithm for learning the RAA and RR values which bootstraps concurrently solved decompositions for coupling on-policy PPO roll-outs. Notably, this allows one to automatically learn to satisfy dual-objective tasks, for example, in the RAA, the F16 learns to fly into the desired airspace without crashing afterward (Figure 1, top middle-left), and in the RR case, the Hopper learns to jump into a target without diving so it may then achieve the second target (Figure 1, bottom left). The RAA and RR problems are distinct from both standard sum-of-reward values and the simpler HJ-RL formulations, providing new perspectives and performant tools for constrained decision-making.
Our contributions include:
  1. 1.
    We introduce novel value functions corresponding to the RAA and RR problems.
  2. 2.
    We prove that these value functions and their optimal policies can be decomposed into reach, avoid, and reach-avoid value functions (Theorems 1 and 2).
  3. 3.
    We demonstrate the nature of the RAA and RR values and their optimal policies in a simple grid-world example with deep $Q$-learning (DQN) (Figure 2).
  4. 4.
    We propose DOHJ-PPO to solve these value functions, which bootstraps concurrently solved decompositions for effectively coupling the on-policy rollouts (Section 7.2).
  5. 5.
    In continuous control tasks, we showcase that with little to no tuning, DOHJ-PPO is more successful, safer and faster than Lagrangian and existing HJ-RL baselines (Figure 4).
2 Related Works
This work involves aspects of safety (e.g. hazard avoidance), liveness (e.g. goal reaching), and balancing competing objectives. We summarize the relevant related works here.
Constrained and Multi-Objective RL. Constrained Markov decision processes (CMPDs) maximize the expected sum of discounted rewards subject to an expected sum of discounted costs, or an instantaneous safety violation function remaining below a set threshold Altman (2021); Achiam et al. (2017a); Wachi and Sui (2020). CMDPs are an effective way to incorporate state constraints into RL problems, and the efficient and accurate solution of the underlying optimization problem has been extensively researched, first by Lagrangian methods and later by an array of more sophisticated techniques Stooke et al. (2020); Li et al. (2024); Chen et al. (2021); Miryoosefi and Jin (2021); Yang et al. (2020). Multi-objective RL is an approach to designing policies that obtain Pareto-optimal expected sums of discounted vector-valued rewards Wiering et al. (2014); Van and Nowé (2014); Cai et al. (2023), including by deep-Q and other deep learning techniques Mossalam et al. (2016); Abels et al. (2019); Yang et al. (2019). By contrast, this work explicitly balances rewards and penalties in a way that does not require specifying a Lagrange multiplier or similar hyperparameter. Moreover, our work treats goal-reaching and hazard-avoidance as hard constraints, and the learned value function has a direct interpretation in terms of the constraint satisfaction.
Goal-Conditioned RL (GCRL). GCRL simultaneously learns optimal policies for a range of different (but typically related) tasks Liu et al. (2022); Plappert et al. (2018); Ren et al. (2019); Ma et al. (2022); Campero et al. (2020); Trott et al. (2019); Eysenbach et al. (2022); Ma et al. (2022); Campero et al. (2020). In GCRL, states are augmented with information on the current goal. While these goals are in their simplest form mostly independent, some work extends GCRL to more sophisticated composite tasks Chane-Sane et al. (2021). Our work primarily focuses on composing specific learned tasks rather than learning general tasks simultaneously.
Linear Temporal Logic (LTL), Automatic State Augmentation, Automatons, and Generalized Objective Functions, . Many works have been explored that merge LTL and RL, canonically focused on Non-Markovian Reward Decision Processes (NMRDPs) Bacchus et al. (1996). Here, the reward gained at each time step may depend on the previous state history. Many of these works convert these NMRDPs to MDPs via state augmentation Bacchus et al. (1997); Thiebaux et al. (2006); Camacho et al. (2021); Icarte et al. (2018); Camacho et al. (2019). Often the augmented states are taken to be products between an ordinary state and an automaton state, where the automaton is used to determine "where" in the LTL specification an agent currently is. Other works using RL for LTL tasks involve MDP verification Brázdil et al. (2014), hybrid systems theory Cohen et al. (2023), GCRL with complex LTL tasks Qiu et al. (2023), almost-sure objective satisfaction Sadigh et al. (2014), incorporating (un)timed specifications Hamilton et al. (2022), and using truncated LTL Li et al. (2017). While the problems we attempt to solve (e.g. reaching multiple goals) can be thought of as specific instantiations of LTL specifications, our approach to solving these problems is fundamentally different from those in this line of work. Our state augmentation and subsequent decomposition of the problem are performed in a specific manner to leverage new HJ-based methods on the subproblems. Through our specific choice of state augmentation, we still prove that we can achieve an optimal policy in theory (and approximately so in practice) despite the non-NMRDP setup. There is also significant literature on generalized objective functions in RL Wang et al. (2020); Cui and Yu (2023); Tang et al. (2025), but these works are either not able to or are not tailored to simultaneously handle multiple rewards/penalties in the context of safe optimal control, which is where our decompositional approach becomes useful. On the other hand, works that do try to handle multiple rewards and penalties (including by decomposition) still use discounted-sum-of-rewards objectives van Seijen et al. (2017); Pitis (2023); Lin et al. (2020).
Hamilton-Jacobi (HJ) Methods. HJ is a dynamic programming-based framework for solving reach, avoid, and reach-avoid tasks Mitchell et al. (2005); Fisac et al. (2015). The value functions used in HJ have the advantage of directly specifying desired behavior, so that a positive value corresponds to task achievement and a negative value corresponds to task failure. Recent works use RL to find corresponding optimal policies by leveraging the unconventional Bellman updates associated with these value functions So et al. (2024); Hsu et al. (2021); Fisac et al. (2019). We build on these works by extending these advancements to more complex tasks, superficially mirroring the progression from MDPs to NMRDPs in the LTL-RL literature. Additional works merge HJ and RL, but do not concern themselves with such composite tasks Ganai et al. (2023); Yu et al. (2022a); Zhu et al. (2024).
3 Problem Definition
@@
Consider a Markov decision process (MDP) $\mathcal{M}=\langle\mathcal{{S}},\mathcal{{A}},f\rangle$ consisting of finite state and action spaces $\mathcal{{S}}$ and $\mathcal{{A}}$, and unknown discrete dynamics $f$ that define the deterministic transition $s_{t+1}=f(s_{t},a_{t})$. Let an agent interact with the MDP by selecting an action with policy $\pi:\mathcal{{S}}\to\mathcal{{A}}$ to yield a state trajectory $s_{t}^{\pi}$, i.e. $s_{t+1}^{\pi}=f\left(s_{t}^{\pi},\pi\left(s_{t}^{\pi}\right)\right).$
@@
In this work, we consider the Reach-Always-Avoid (RAA) and Reach-Reach (RR) problems, which both involve the composition of two objectives, which are each specified in terms of the best reward and worst penalty encountered over time. In the RAA problem, let $r,p:\mathcal{{S}}\to\mathbb{R}$ represent a reward to be maximized and a penalty to be minimized. We will let q=pq=-p for mathematical convenience, but for conceptual ease we recommend the reader think of trying to minimize the largest-over-time penalty $p$ rather than maximize the smallest-over-time $q$. In the RR problem, let $r_{1},r_{2}:\mathcal{{S}}\to\mathbb{R}$ be two distinct rewards to be maximized. The agent’s overall objective is to maximize the worst-case outcome between the best-over-time reward and worst-over-time penalty (in RAA) and the two best-over-time rewards (in RR), i.e.
$(\textsc{RAA})$ $\left\{\begin{array}[]{cl}{\operatorname{maximize}\text{ (w.r.t. }\pi)}&\min\big\{\max_{t}r(s_{t}^{\pi}),\>\min_{t}q(s_{t}^{\pi})\big\}\\ \text{ s.t. }&s_{t+1}^{\pi}=f\left(s_{t}^{\pi},\pi\left(s_{t}^{\pi}\right)\right),\\ &s_{0}^{\pi}=s,\end{array}\right$.
$(\textsc{RR})$ $\left\{\begin{array}[]{cl}{\operatorname{maximize}\text{ (w.r.t. }\pi)}&\min\big\{\max_{t}r_{1}(s_{t}^{\pi}),\>\max_{t}r_{2}(s_{t}^{\pi})\big\}\\ \text{ s.t. }&s_{t+1}^{\pi}=f\left(s_{t}^{\pi},\pi\left(s_{t}^{\pi}\right)\right),\\ &s_{0}^{\pi}=s.\end{array}\right$.
@@
As the names suggest, these optimization problems are inspired by — but not limited to — tasks involving goal reaching and hazard avoidance. More specifically, the RAA problem is motivated by a task in which an agent wishes to both reach a goal $\mathcal{G}$ and perennially avoid a hazard $\mathcal{H}$ (even after it reaches the goal). The RR problem is motivated by a task in which an agent wishes to reach two goals, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, in either order. While these problems are thematically distinct, they are mathematically complementary (differing by a single $\max$$\min$ operation), and hence we tackle them together.
Figure 2: DQN Grid-World Demonstration of the RAA & RR Problems. We compare our novel formulations with previous HJ-RL formulations (RA & R) in a simple grid-world problem with DQN. The zero-level sets of $q$ (hazards) are highlighted in red, those of $r$ (goals) in blue, and trajectories in black (starting at the dot). In both models, the agents actions are limited to {left, right, straight} and the system flows upwards over time.
The values for any policy in these problems then take the forms $V_{\textup{RAA}}^{\pi}$ and $V_{\textup{RR}}^{\pi}$,
$V_{\textup{RAA}}^{\pi}(s)=\min\left\{\max_{t}r(s_{t}^{\pi}),\>\min_{t}q(s_{t}^{\pi})\right\}\quad\text{and}\quad V_{\textup{RR}}^{\pi}(s)=\min\left\{\max_{t}r_{1}(s_{t}^{\pi}),\>\max_{t}r_{2}(s_{t}^{\pi})\right\}.$
One may observe that these values are fundamentally different from the infinite-sum value commonly employed in RL Sutton and Barto (2018), and do not accrue over the trajectory but, rather, are determined by certain points. Moreover, while each return considers two objectives, these objectives are combined in worst-case fashion to ensure dual-satisfaction. Although many of the works discussed in the previous section approach related tasks (e.g. goal reaching and hazard avoidance) via traditional sum-of-discounted-rewards formulations, these novel value functions have a more direct interpretation in the following sense: if $r$ is positive (only) within $\mathcal{G}$ and $q$ is positive (only) inside $\mathcal{H}$, $V_{\textup{RAA}}^{\pi}(s)$ will be positive if and only if the RAA task will be accomplished by the policy $\pi$. Similarly if $r_{1}$ and $r_{2}$ are positive within $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively, $V_{\textup{RR}}^{\pi}(s)$ will be positive if and only if the RR task will be accomplished by the policy $\pi$.
4 Reachability and Avoidability in RL
The reach $V_{\textup{R}}^{\pi}$, avoid $V_{\textup{A}}^{\pi}$, and reach-avoid $V_{\textup{RA}}^{\pi}$ values, respectively defined by
$V_{\textup{R}}^{\pi}(s)=\max_{t}r(s_{t}^{\pi}),\quad V_{\textup{A}}^{\pi}(s)=\min_{t}q(s_{t}^{\pi}),\quad V_{\textup{RA}}^{\pi}(s)=\max_{t}\min\left\{r(s_{t}^{\pi}),\min_{\tau\leq t}q(s_{\tau}^{\pi})\right\},$
@@ 4 Reachability and Avoidability in RL
have been previously studied Fisac et al. (2019) leading to the derivation of special Bellman equations. To put these value functions in context, assume the goal $\mathcal{G}$ is the set of states for which $r(s)$ is positive and the hazard $\mathcal{H}$ is the set of states for which $q(s)$ is non-positive. See Figure 2 for a simple grid-world demonstration comparing the RAA and RR values with the previously existing RA and R values. Then $V_{\textup{R}}^{\pi}$, $V_{\textup{A}}^{\pi}$, and $V_{\textup{RA}}^{\pi}$ are positive if and only if $\pi$ causes the agent to eventually reach $\mathcal{G}$, to always avoid $\mathcal{H}$, and to reach $\mathcal{G}$ without hitting $\mathcal{H}$ prior to the reach time, respectively. The Reach-Avoid Bellman Equation (RABE), for example, takes the form Hsu et al. (2021)
@@
and is associated with optimal policy $\pi_{\text{RA}}^{*}(s)$ (without the need for state augmentation, see the appendix). This formulation does not naturally induce a contraction, but may be discounted to induce contraction by defining $V_{\textup{RA}}^{\gamma}(z)$ implicitly via
$V_{\textup{RA}}^{\gamma}(s)=(1-\gamma)\min\{r(s),q(s)\}+\gamma\min\left\{\max\left\{\max_{a\in\mathcal{{A}}}V_{\textup{RA}}^{\gamma}\left(f(s,a)\right),r(s)\right\},q(s)\right\},$
for each $\gamma\in[0,1)$. A fundamental result (Proposition 3 in Hsu et al. (2021)) is that
$\lim_{\gamma\to 1}V_{\textup{RA}}^{\gamma}(s)=V_{\textup{RA}}(s).$
@@
These prior value functions and corresponding Bellman equations have proven powerful for these simple reach/avoid/reach-avoid problem formulations. In this work, we generalize the aforementioned results to the broader class involving $V_{\textup{RAA}}$ (assure no penalty after the reward threshold is achieved) and $V_{\textup{RR}}$ (achieve multiple rewards optimally). Through this generalization, we are able to train an agent to accomplish more complex tasks with noteworthy performance.
5 The need for augmenting states with historical information
@@
We here discuss a small but important detail regarding the problem formulation. The value functions we introduce may appear similar to the simpler HJ-RL value functions discussed in the previous section; however, in these new formulations the goal of choosing a policy $\pi:\mathcal{{S}}\to\mathcal{{A}}$ is inherently flawed without state augmentation. In considering multiple objectives over an infinite horizon, situations arise in which the optimal action depends on more than the current state, but rather the history the trajectory. This complication is not unique to our problem formulation, but also occurs for NMDPs (see the Related Works section). To those unfamiliar with NMDPs, this at first may seem like a paradox as the MDP is by definition Markov, but the problem occurs not due to the state-transition dynamics but the nature of the reward. An example clarifying the issue is shown in Figure 3.
@@
Figure 3: Examples where a Non-Augmented Policy is Flawed. In both MDPs, consider an agent with no memory. (Left) For a deterministic policy based on the current state, the agent can only achieve one target ($R$), as this policy must associate the middle state with either of the two possible actions. (Right) The RAA case is slightly more complex. Assume the robot will make sure to avoid the fire at all costs (which is easily done from the current state). It would also prefer to not encounter the cone hazard, but will do so if needed to achieve the target. From its current state the robot cannot determine whether to pursue the target by crossing the cone or move to the right. The correct decision depends on state history, specifically on whether the robot has already reached the target state or not (e.g. imagine the initial state is on the target state).
@@
To allow the agent to use relevant aspects of its history, we will henceforth consider an augmentation of the MDP with auxiliary variables. A theoretical result in the next section states that this choice of augmentation is sufficient in that no additional information will be able to improve performance under the optimal policy. Note that the state augmentation is needed because of the use of HJR-style optimization objectives (rather than discounted sum-of-rewards). The point of the state augmentation is not to make the rewards and penalties Markovian (indeed, they are already Markovian as they are deterministic functions of the current state).
5.1 Augmentation of the RAA Problem